# Constraining $A\to ZH$ with $H\to t\bar{t}$ in the Low-Mass Region

### v2

@@ Abstract
The decay $A\to ZH$ is a characteristic signal of two-Higgs-doublet models (2HDMs), where $A$ and $H$ lie primarily within the same $SU(2)_{L}$ multiplet, leading to a coupling of order $g_{2}$ to the $Z$ boson. The subsequent decay $H\to tt^{(*)}$ is particularly promising, as it gives rise to distinct final states involving multiple leptons and $b$-jets. The required splitting between $m_{A}$ and $m_{H}$ can naturally occur near the electroweak scale while being consistent with perturbative unitarity. Whereas dedicated ATLAS and CMS searches focused on the region with both top-quarks on-shell, we cover lower masses where one top quark is off-shell by recasting Standard Model $t\bar{t}Z$ measurements of ATLAS and CMS. The obtained limits on $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})$ are between 0.12 pb and 0.62 pb. Interestingly, we observe these stringent limits despite a preference (up to $2.5\sigma$) for a non-zero new physics signal, most pronounced around for $m_{A}\approx 450$–460 GeV and $m_{H}\approx 290$ GeV, with a best-fit value of $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})\approx 0.3$ pb. This cross section can be accommodated within a top-philic 2HDM for a top-Yukawa coupling of the second Higgs doublet of $\mu_{t}\gtrsim 0.16$.
@@ I Introduction
The search for new Higgs bosons is a primary goal of the LHC program. Of particular interest are the cascade decays of heavy Higgs states (see Refs. [30, 31] for an overview), such as the decay of a pseudoscalar Higgs boson $A$ into a $Z$ boson and a new scalar $H$ ($A\to ZH$). [^1] While we considered $A\to ZH$, one can also study $H\to AZ$ which can be distinguished from the former via spin correlations [32]. Notably, this chain decay has been identified as a ‘smoking gun’ signature for a strong first-order electroweak phase transition, which could explain the origin of the matter-antimatter asymmetry in the Universe [33, 34, 35]. Furthermore, owing to their clean experimental signatures, these processes provide powerful probes of extended Higgs sectors. In particular, the decay $H\to t\bar{t}$ yields distinctive multi-lepton final states with $b$-jets, and has been searched for above the top threshold by ATLAS [36] and CMS [37]. [^2] ATLAS considers $m_{H}>350$ GeV and CMS $m_{H}>330$ GeV. Since CMS does not mention a $t\bar{t}$ threshold, it seems that they used the $\overline{\text{MS}}$ mass for the top quark, whereas for the decay width the on-shell mass would be more appropriate. We therefore show the ATLAS search region in our plots.
@@ II Analyses of $t\bar{t}Z$ differential distributions
Figure 3: 95% CL upper limit (left) and best-fit (right) value of $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})$ in units of pb in the $m_{A}$$m_{H}$ plane with $m_{A}-m_{H}\geq 100$ GeV, obtained by combining the CMS and ATLAS analyses. The color bar in the left (right) plot indicates the 95% CL upper limit for the cross section (the preference for a non-zero NP signal in units of standard deviation). An extended region with $m_{A}<800\,\mathrm{GeV}$ is provided in Fig. 8 in the Appendix.
For the reconstruction and selection of the leptons (electrons and muons) and jets (including $b$-tagged jets), we closely follow the CMS and ATLAS analyses; jets are clustered with the anti-$k_{T}$ algorithm [47] implemented in FastJet 3.3.4 [48]. Events are required to contain at least three leptons, including a same-flavor opposite-sign pair with an invariant mass in the nominal $Z$ boson window, and a third lepton consistent with a $W$ boson decay. All additional selection requirements—jet and $b$-tagged jet multiplicities and kinematic thresholds on leptons and jets—are applied to reproduce the event selections used in the CMS and ATLAS signal regions.
We then simulate our new physics (NP) signal $pp\to A\to ZH$ with $H\to t\bar{t}$ using the same setup. To have an on-shell $Z$ boson, we require a mass splitting $m_{A}-m_{H}\geq 100$ GeV. Our scan covers $m_{H}$ from 260 GeV to 400 GeV (such that one top quark and the $W$ boson originating from the off-shell top quark are on-shell) and $m_{A}$ from 360 GeV to 800 GeV. The resulting differential distributions for the benchmark point $(m_{A},m_{H})=(460,290)$ GeV and $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})=0.3$ pb are shown in the Appendix (Fig. 6 and Fig. 7).
@@ III Results and Interpretation
represents the expected cross section in bin $i$, with SM and NP contributions weighted by the corresponding fit parameters; $\mu_{\rm NP}$ is identified with $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})$. Correlations among the differential observables are obtained from our SM simulation, following Ref. [49].
@@
Figure 4: Left: Regions in the $m_{A}$$m_{H}$ plane for the top-philic 2HDM, assuming $\sin\alpha=0$, that are excluded by our analysis for different values of the top Yukawa rescaling parameter $\mu_{t}$. Right: Corresponding exclusion in the aligned 2HDM realization with $\xi_{u}\neq 0$, i.e., both top and charm couplings.
@@ III.1 Interpretation
For the best-fit region around $m_{A}\simeq 450$–460 GeV and $m_{H}\simeq 290$  GeV (see right plot in Fig. 3), the preference for a non-vanishing NP contribution can be accommodated for $\mu_{t}\gtrsim 0.16$ and $\sin\alpha\sim 0$; see Fig. 5. The $1\sigma$ and $2\sigma$ regions (shaded in blue) are shown together with the exclusions from the CMS and ATLAS searches for $A\to t\bar{t}$ [58, 59]. The corresponding best-fit line (purple) for the aligned realization is also displayed. The regions above the red and green lines are excluded by the ATLAS and CMS searches for $A\to t\bar{t}$ [58, 59], respectively.
@@ IV Conclusions and Outlook
Interpreting the results within the top-philic 2HDM, novel constraints are derived. Furthermore, the observed preference for a non-zero NP signal can be accommodated in the top-philic 2HDM for $\mu_{t}\gtrsim 0.16$. Future high-luminosity LHC measurements of differential $t\bar{t}Z$ production and dedicated searches for $A\to ZH$ with $H\to t\bar{t}$ will allow this region to be probed with greater precision, providing a direct test of the top-philic Higgs scenario like the aligned 2HDM with $\xi_{d,\ell}\approx 0$ but $\xi_{u}\neq 0$. In general, our analysis shows the importance and feasibility of searching for Higgs signals with off-shell top quarks in the final state.
@@ Acknowledgements.
SA is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 396021762 – TRR 257. AC is supported by a professorship grant of the Swiss National Science Foundation (Grant No. PP00P21_76884). SPM and BM acknowledge the support of the Research Office of the University of the Witwatersrand. BM further acknowledges support from the South African Department of Science and Innovation through the SA-CERN program, and the National Research Foundation. We thank Thomas Biekötter for pointing out the ATLAS search for $A\to ZH\to\ell\ell bb$ [62], which excludes a sizable part of the type-I 2HDM parameter space. *
Appendix A
@@
Finally, Fig. 8 displays the extended $m_{A}$$m_{H}$ plane with the resulting 95% CL upper limits on $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})$, extending the results of Fig. 3.
Figure 6: $t\bar{t}Z+tWZ$ differential cross sections for various observables measured by CMS. The error bars indicate the total experimental error, while the shaded blue area corresponds to the theory uncertainty. The NP contributions are shown for $(m_{A},m_{H})=(460,290)$ GeV and $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})=0.3$ pb, corresponding to the combined (ATLAS and CMS) best-fit point.
@@
Figure 8: Extended 95% CL upper limits on $\sigma(A\to ZH)\times\mathrm{Br}(H\to t\bar{t})$ in units of pb in the $m_{A}$$m_{H}$ plane with $m_{A}-m_{H}\geq 100~\mathrm{GeV}$. This plot extends the results shown in Fig. 3 to a wider mass range.

### v3

@@ Abstract
The decay $A\to ZH$ is a characteristic signal of two-Higgs-doublet models (2HDMs), where $A$ and $H$ lie primarily within the same $SU(2)_{L}$ multiplet, leading to a coupling of order $g_{2}$ to the $Z$ boson. The subsequent decay $H\to tt^{(*)}$ is particularly promising, as it gives rise to distinct final states involving multiple leptons and $b$-jets. The required splitting between $m_{A}$ and $m_{H}$ can naturally occur near the electroweak scale while being consistent with perturbative unitarity. Whereas dedicated ATLAS and CMS searches focused on the region with both top-quarks on-shell, we cover lower masses where one top quark is off-shell by recasting Standard Model $t\bar{t}Z$ measurements of ATLAS and CMS. The obtained limits on $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})$ are between 0.12 pb and 0.62 pb. Interestingly, we observe these stringent limits despite a preference (up to $2.5\sigma$) for a non-zero new physics signal, most pronounced around for $m_{A}\approx 450$–460 GeV and $m_{H}\approx 290$ GeV, with a best-fit value of $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})\approx 0.3$ pb. This cross section can be accommodated within a top-philic 2HDM for a top-Yukawa coupling of the second Higgs doublet of $0.16\lesssim\mu_{t}\lesssim 0.33$.
@@ I Introduction
The search for new Higgs bosons is a primary goal of the LHC program. Of particular interest are the cascade decays of heavy Higgs states (see Refs. [30, 31] for an overview), such as the decay of a pseudoscalar Higgs boson $A$ into a $Z$ boson and a new scalar $H$ ($A\to ZH$).Notably, this chain decay has been identified as a ‘smoking gun’ signature for a strong first-order electroweak phase transition, which could explain the origin of the matter-antimatter asymmetry in the Universe [32, 33, 34]. While we focus on the decay $A\to ZH$, a strong first-order electroweak phase transition can also be realized when the channel $H\to ZA$ is kinematically open, even though parameter scans often favor the former [35, 36]. From a collider perspective, both decay chains lead to very similar final states when followed by $A,H\to t\bar{t}$. [^1] The two processes can nevertheless be distinguished using spin correlations [37]. Furthermore, owing to their clean experimental signatures, these processes provide powerful probes of extended Higgs sectors. In particular, the decay $H\to t\bar{t}$ yields distinctive multi-lepton final states with $b$-jets, and has been searched for above the top threshold by ATLAS [38] and CMS [39]. [^2] ATLAS considers $m_{H}>350$ GeV and CMS $m_{H}>330$ GeV. Since CMS does not mention a $t\bar{t}$ threshold, it seems that they used the $\overline{\text{MS}}$ mass for the top quark, whereas for the decay width the on-shell mass would be more appropriate. We therefore show the ATLAS search region in our plots.
@@ II Analyses of $t\bar{t}Z$ differential distributions
Figure 3: 95% CL upper limit (left) and best-fit (right) value of $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})$ in units of pb in the $m_{A}$$m_{H}$ plane with $m_{A}-m_{H}\geq 100$ GeV, obtained by combining the CMS and ATLAS analyses. The color bar in the left (right) plot indicates the 95% CL upper limit for the cross section (the preference for a non-zero NP signal in units of standard deviation). An extended region with $m_{A}<800\,\mathrm{GeV}$ is provided in the Supplemental Material.
For the reconstruction and selection of the leptons (electrons and muons) and jets (including $b$-tagged jets), we closely follow the CMS and ATLAS analyses; jets are clustered with the anti-$k_{T}$ algorithm [49] implemented in FastJet 3.3.4 [50]. Events are required to contain at least three leptons, including a same-flavor opposite-sign pair with an invariant mass in the nominal $Z$ boson window, and a third lepton consistent with a $W$ boson decay. All additional selection requirements—jet and $b$-tagged jet multiplicities and kinematic thresholds on leptons and jets—are applied to reproduce the event selections used in the CMS and ATLAS signal regions, and are summarized explicitly in the Supplemental Material.
We then simulate our new physics (NP) signal $pp\to A\to ZH$ with $H\to t\bar{t}$ using the same setup. To have an on-shell $Z$ boson, we require a mass splitting $m_{A}-m_{H}\geq 100$ GeV. Our scan covers $m_{H}$ from 260 GeV to 400 GeV (such that one top quark and the $W$ boson originating from the off-shell top quark are on-shell) and $m_{A}$ from 360 GeV to 800 GeV. The resulting differential distributions for the benchmark point $(m_{A},m_{H})=(460,290)$ GeV and $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})=0.3$ pb are shown in in the Supplemental Material.
@@ III Results and Interpretation
represents the expected cross section in bin $i$, with SM and NP contributions weighted by the corresponding fit parameters; $\mu_{\rm NP}$ is identified with $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})$. Interference effects between the SM and NP contributions are neglected; given the narrow width of $A$ in the parameter space considered, these effects are expected to be subleading compared to current experimental uncertainties. Correlations among the differential observables are obtained from our SM simulation, following Ref. [51].
@@
Figure 4: Left: Regions in the $m_{A}$$m_{H}$ plane for the top-philic 2HDM, assuming $\sin\alpha=0$, that are excluded by our analysis for different values of the top Yukawa rescaling parameter $\mu_{t}$. Right: Corresponding exclusion in the aligned 2HDM realization with $\xi_{u}\neq 0$, i.e., both top and charm couplings. An extended version of this figure covering a larger $m_{A}$ range up to 800 GeV is provided in the Supplemental Material.
@@ III.1 Interpretation
For the best-fit region around $m_{A}\simeq 450$–460 GeV and $m_{H}\simeq 290$  GeV (see right plot in Fig. 3), the preference for a non-vanishing NP contribution can be accommodated for $0.16\lesssim\mu_{t}\lesssim 0.33$ and $\sin\alpha\sim 0$; see Fig. 5. The $1\sigma$ and $2\sigma$ regions (shaded in blue) are shown together with the exclusions from the CMS and ATLAS searches for $A\to t\bar{t}$ [60, 61]. The corresponding best-fit line (purple) for the aligned realization is also displayed. The regions above the red and green lines are excluded by the ATLAS and CMS searches for $A\to t\bar{t}$ [60, 61], respectively.
@@ IV Conclusions and Outlook
Interpreting the results within the top-philic 2HDM, novel constraints are derived. Furthermore, the observed preference for a non-zero NP signal can be accommodated in the top-philic 2HDM for $0.16\lesssim\mu_{t}\lesssim 0.33$. Future high-luminosity LHC measurements of differential $t\bar{t}Z$ production and dedicated searches for $A\to ZH$ with $H\to t\bar{t}$ will allow this region to be probed with greater precision, providing a direct test of the top-philic Higgs scenario like the aligned 2HDM with $\xi_{d,\ell}\approx 0$ but $\xi_{u}\neq 0$. For completeness, we note that in 2HDMs with natural flavour conservation, the top Yukawa coupling is suppressed away from the low-$\tan\beta$ regime. As a consequence, in the mass region considered here, searches for $pp\to A\to ZH$ with $H\to b\bar{b}$ [64] typically provide stronger constraints than the $H\to t\bar{t}$ channel. Our results are therefore complementary and mainly relevant for scenarios with enhanced top couplings. In general, our analysis shows the importance and feasibility of searching for Higgs signals with off-shell top quarks in the final state.
@@ Acknowledgements.
SA is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 396021762 – TRR 257. AC is supported by a professorship grant of the Swiss National Science Foundation (Grant No. PP00P21_76884). SPM and BM acknowledge the support of the Research Office of the University of the Witwatersrand. BM further acknowledges support from the South African Department of Science and Innovation through the SA-CERN program, and the National Research Foundation. We thank Thomas Biekötter for pointing out the ATLAS search for $A\to ZH\to\ell\ell bb$ [64], which excludes a sizable part of the type-I 2HDM parameter space.
Supplemental Material
S1 CMS and ATLAS Event Selections
We now summarize the event selections applied in our recast to reproduce the fiducial $t\bar{t}Z$ regions used by CMS and ATLAS for their differential measurements.
To reproduce the CMS $t\bar{t}Z$ signal regions [41], we require events with at least three charged leptons ($e,\mu$). Leptons are required to satisfy $|\eta|<2.5$ for electrons and $|\eta|<2.4$ for muons, with transverse-momentum thresholds of $p_{T}>25$, $15$, and $10$ GeV for the three leading leptons. A $Z$-boson candidate is reconstructed from a same-flavor opposite-sign lepton pair with invariant mass $70<m_{\ell\ell}<110$ GeV, while the third lepton is interpreted as originating from the $W$ boson decay. Jets are reconstructed using the anti-$k_{T}$ algorithm with $R=0.4$, requiring $p_{T}^{j}>25$ GeV and $|\eta_{j}|<5.0$, and are required to be separated from leptons by $\Delta R(j,\ell)>0.4$. Events are required to contain at least two jets, of which at least one must be $b$-tagged.
To reproduce the ATLAS $t\bar{t}Z$ fiducial selections used for the differential measurements, we follow the particle-level definitions given in Ref. [40]. In the trilepton channel, we require exactly three charged leptons ($e,\mu$) with $p_{T}>27,20,15$ GeV and total charge $\pm 1$, at least one opposite-sign same-flavor (OSSF) lepton pair with $|m_{\ell\ell}-m_{Z}|<10$ GeV, and $m_{\mathrm{OSSF}}>10$ GeV for all OSSF combinations. We further require at least three jets with $p_{T}^{j}>25$ GeV and $|\eta_{j}|<2.5$, and at least one $b$-tagged jet. In the tetralepton channel, we require exactly four leptons with $p_{T}>27,7,7,7$ GeV and total charge zero, at least one OSSF lepton pair with $|m_{\ell\ell}-m_{Z}|<20$ GeV, and $m_{\mathrm{OSSF}}>10$ GeV for all OSSF combinations, together with at least two jets with $p_{T}^{j}>25$ GeV and $|\eta_{j}|<2.5$, and at least one $b$-tagged jet.
Figure 6: $t\bar{t}Z+tWZ$ differential cross sections for various observables measured by CMS. The error bars indicate the total experimental error, while the shaded blue area corresponds to the theory uncertainty on the SM prediction as provided by the CMS differential measurement, including contributions from scale variations, PDFs, and parton-shower modelling. The NP contributions are shown for $(m_{A},m_{H})=(460,290)$ GeV and $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})=0.3$ pb, corresponding to the combined (ATLAS and CMS) best-fit point. The corresponding couplings in the top-philic 2HDM depend on $\alpha$ and $\mu_{t}$, with the region associated with this benchmark shown in Fig. 5 of the main text.
S2 Differential Cross Sections for CMS and ATLAS Observables
@@
S3 Extended limits in the $m_{A}$$m_{H}$ plane
Finally, Fig. 8 displays the extended $m_{A}$$m_{H}$ plane with the resulting 95% CL upper limits on $\sigma(A\to ZH)\times{\rm Br}(H\to t\bar{t})$, extending the results of Fig. 3 of the main text. In addition, Fig. 9 displays the corresponding extension of the 2HDM interpretation, illustrating the largest excluded pseudoscalar masses for large values of $\mu_{t}$.
Figure 8: Extended 95% CL upper limits on $\sigma(A\to ZH)\times\mathrm{Br}(H\to t\bar{t})$ in units of pb in the $m_{A}$$m_{H}$ plane with $m_{A}-m_{H}\geq 100~\mathrm{GeV}$. This plot extends the results shown in Fig. 3 of the main text to a wider mass range.
Figure 9: Same as Fig. 4, but extended to $m_{A}<800$ GeV to illustrate the largest excluded masses for large $\mu_{t}$.